OpenStudy (kathydimieri12):

Which expression is equivalent to 2(9 + 7)? A. 18 + 14 B. 18 + 7 C. 11 + 9 D. 11 + 7

OpenStudy (mathewr.):

18 + 14

OpenStudy (kathydimieri12):

Why?

OpenStudy (anonymous):

just distribute \[\large \sf a(x+y)=ax+ay\]

OpenStudy (anonymous):

So \[\large \sf 2(9+7)=2 \times 9 + 2 \times 7\]

OpenStudy (mathewr.):

because when you distribute its like 9 x 2 and 2 x 7 18 + 14

OpenStudy (ihelpyoulearn):

This would be 18+14. I know this because you would take 2 and times it by both 9 and 7 which are in the parenthesis. So 2x9=18 and 2x7=14. Then you add both 18 and 14 and get 32.

OpenStudy (anonymous):

A they both r equal to 32

OpenStudy (kathydimieri12):

o thank you all

OpenStudy (ihelpyoulearn):

No problem :)

OpenStudy (kathydimieri12):

:)

OpenStudy (anonymous):

np ;)